OpenStudy (anonymous):

Which of the following is a correct equation for the line passing through the point (-2,1) and having slope m = 1/2? Check all that apply. A.y=-2x+1/2 B.y=1/2x+2 C.y-1=1/2(x+2) D.x-2y=-4

OpenStudy (anonymous):

y=(x/2)+2

OpenStudy (anonymous):

@LewisSparrow96

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

what thats not one the answers

OpenStudy (anonymous):

@Princer_Jones

OpenStudy (anonymous):

From y=(x/2)+2 we have 2y=x+4 and hence x-2y=-4 That is option D

OpenStudy (anonymous):

@LewisSparrow96

OpenStudy (anonymous):

Hei please check it now @LewisSparrow96